The molar fraction of solute is 0.35

a. Calculate the mole fraction of solvent, H2O.

b. What is the mole of solute if the mole of solution is 4.11mole?

c. Find the mass of the solvent.



1
Expert's answer
2021-10-10T10:20:08-0400

Mole fraction (Xsolute) ="\frac{moles of solute }{ total moles of. Solution}"


(a) Mole fraction of water =0.2


(b) Applying the Formula

0.35="\frac{moles of solute}{4.11}"


Moles of solute =0.35×4.11=1.4385moles


(c) molality = "\frac{moles of solute}{ mass of solvent in kilograms}"

0.2M=1.4385

=7.2kg